OpenStudy (anonymous):

−2x + 7y = 13 3x + 2y = −7

OpenStudy (anonymous):

x=-3, y=1

OpenStudy (anonymous):

please show work, thanks

OpenStudy (anonymous):

Show me yours and I'll help you if you get stuck.

OpenStudy (anonymous):

I have never done a problem thats why I am asking to show work, my teacher says to look in our book but I can't find it

OpenStudy (anonymous):

What method do they ask you to use to solve it. There are several ways to do it.

OpenStudy (anonymous):

if you want to eliminate the x-variabls through a linear combination of the two equation, you would multiply the first equation by 2 and the second equation by 3. that would make the coefficient of the x-variables opposites -6x + 21y =39 6x + 4y =-14 ========== 25y=25 y=1 back substitute into any equation above ego get x; i choose the original equation 2 3x +2(1) = -7 3x = -9 x = -3 the solution is x=-3 and y=1, or as an ordered pair (-3, 1)